Three resistances of 12, 16, and 20 ohms are connected in parallel. What resistance must be connected in series with this combination to give a total resistance of 25 ohms?


1
Expert's answer
2021-01-15T13:46:49-0500

Let's first find the total resistance of three resistances connected in parallel:


"\dfrac{1}{R_{tot}}=\dfrac{1}{R_{1}}+\dfrac{1}{R_{2}}+\dfrac{1}{R_{3}},""R_{tot}=\dfrac{1}{\dfrac{1}{R_{1}}+\dfrac{1}{R_{2}}+\dfrac{1}{R_{3}}},""R_{tot}=\dfrac{1}{\dfrac{1}{12\ \Omega}+\dfrac{1}{16\ \Omega}+\dfrac{1}{20\ \Omega}}=5\ \Omega."

Finally, we can find the resistance that must be connected in series with this combination to give a total resistance of 25 ohms:


"R_{tot,parallel}+R=R_{tot,series},""R=R_{tot,series}-R_{tot,parallel}=25\ \Omega-5\ \Omega=20\ \Omega."

Answer:

"R=20\ \Omega."